# Tutorial on amortized optimization

### v4

@@ 3.4.2 Neural acceleration for matrix factorization (NeuralNMF)
where the basis matrix $W\in\mathbb{R}^{m\times r}$ and mixture matrix $H\in\mathbb{R}^{n\times r}$ are elementwise non-negative matrices of rank $r\leq\min(m,n)$. Let $Z=(W,H)$. Taking the norm of the residual of eq. 3.22 leads to the optimization formulation
@@ 3.4.3 HyperAnderson acceleration and deep equilibrium models (HyperDEQ)
Deep equilibrium (DEQ) models [Bai et al., 2019, 2020] investigate implicit layers that parameterize and solve fixed-point computations and have been a flagship for “infinite depth” vision and language models. Given an input $x\in{\mathcal{X}}$, such as an image or language sequence to process, a DEQ model finds a fixed point $y^{\star}(x)$ of $g_{\varphi}(y;x)$ to make a prediction for a task, such as regression or classification. This fixed-point problem can again be interpreted as finding the minimum norm of the residual ${\mathcal{N}}(y;x)\vcentcolon=||y-g_{\varphi}(y;x)||_{2}^{2}$ as
@@ 3.4.4 Comparing NeuralFP and HyperAA
$\tilde{z}^{t+1}$ :$\vcentcolon=z^{t}+{\rm diag}(\rho^{t})^{-1}(v^{t+1}-y^{t})$ (3.29)
$z^{t+1}$ :$\vcentcolon=\Pi\left(\tilde{z}^{t+1}+{\rm diag}(\rho^{t})^{-1}y^{t}\right)$
$y^{t+1}$ :$\vcentcolon=x^{t}+{\rm diag}(\rho)\left(\tilde{z}^{t+1}-z{t+1}\right),$
@@
RLQP proposes to use these OSQP iterates as a semi-amortized model with the iterates $\{s^{t},\rho^{t}\}$. The propose to only parameterize and learn to predict the step size $\rho^{t+1}\vcentcolon=\pi_{\theta}(s^{t})$, with a neural network amortization model $\pi_{\theta}$. They model the process of predicting the optimal $\rho$ as an MDP and define a reward $R_{\rm RLQP}(s,\rho)$ that is $-1$ if the QP is not solved (based on thresholds of the primal and dual residuals) and 00 otherwise, i.e. each episode rolls out the OSQP iterations with a policy predicting the optimal step size. They solve this MDP with TD3 by Fujimoto et al. [2018] to find the parameters $\theta$.
@@ 3.5 Optimal transport
where the coupling $\pi$ is joint distributions over the product of the measures, ${\mathcal{U}}$ is the set of admissible couplings, $c$ is a cost. The dual of eq. 3.30, e.g. in Peyré et al. [2019, Eq. 2.31], can be represented by
@@
is the $c$-transform operation. In Euclidean spaces with the negative inner product cost [Villani, 2009, eq. 5.12], $f$ is a convex function and the $c$-transform operation $f^{c}$ in eq. 3.33 is the standard Legendre-Fenchel transform, also known as the convex conjugate. When the measures $\alpha,\beta$ are discrete, a coupling $\pi$ in the primal (eq. 3.30) can be represented as a matrix, and the potential $f$ in the dual (eq. 3.31) can be represented as a finite-dimensional vector and solved with a linear programming formulation or Sinkhorn iterations [Cuturi, 2013] in the entropic setting.
The following sections overview methods that amortize these optimization problems arising for optimal transport. Section 3.5.1 discusses methods that amortize multiple OT problems and map from the measures and cost to the optimal coupling in eq. 3.30 or duals in eq. 3.31. Section 3.5.2 discusses methods that amortize the $c$-transform operation in eq. 3.33 that arises as a repeatedly-solved subproblem when solving a single OT problem. Section 3.5.3 overviews amortizing a slicing optimization problem that arises when projecting measures down to a single dimension for more computationally efficient solves.
@@ 3.5.1 Amortizing solutions to optimal transport problems
Many computational OT methods focus on computing the solution mapping from the measures and cost to the optimal coupling (eq. 3.30) or duals (eq. 3.31). When repeatedly solving optimal transport problems, this solution mapping is an optimization problem that can be amortized. These methods for predicting the optimal duals of OT problems are also related to Dinitz et al. [2021], Khodak et al. [2022], which predicts the dual solutions to matching problems. They are also related to other heuristic-based initializations for OT problems that do not use amortization such as Thornton and Cuturi [2022].
@@ 3.5.2 Amortized convex conjugates (AmorConj) and $c$-transforms
Most methods for computing the dual in eq. 3.31 between a single pair of measures needs to repeatedly compute the $c$-transform in eq. 3.33 to evaluate the objective value $J$ in eq. 3.32. This transform is typically not a computational bottleneck in discrete settings when the measures have a few thousand points, e.g. in Sinkhorn solvers such as [Cuturi, 2013]. Otherwise in some continuous settings, the conjugate operation may be computationally challenging because it is an optimization problem that needs to be repeatedly solved from scratch to obtain a single Monte-Carlo estimate of $\int_{\mathcal{Y}}f^{c}(y){\rm d}\beta(y)$ to evaluate $J$ in eq. 3.32 once.
Scoping to computing the optimal transport maps between Euclidean measures with the negative inner product cost [Villani, 2009, eq. 5.12], $f$ is a convex function and the $c$-transform operation $f^{c}$ in eq. 3.33 is the standard Legendre-Fenchel transform. Taghvaei and Jalali [2019] discusses a lot of the theoretical foundations underpinning Wasserstein-2 optimal transport and experimentally use a numerical method that solves each conjugate operation from scratch. To alleviate the computational bottleneck of this, many methods using similar theoretical foundations amortize this conjugate operation [Dam et al., 2019, Makkuva et al., 2020, Korotin et al., 2021, Amos, 2023]. The simplest instantiation of this amortization uses a fully-amortized model $\hat{x}_{\theta}(y)$ trained with objective-based learning of the conjugate objective. Amos [2023] further discusses the modeling and loss choices in this setting and also discusses the idea of fine-tuning the amortized prediction with a numerical solver to ensure the dual objective is accurately estimated.
@@ 3.5.3 Amortized Sliced Wasserstein (${\mathcal{A}}$-SW) by Nguyen and Ho [2022]
[…]
where $\theta_{\#}\alpha$ is the push-forward measure of $\mu$ through $T_{\theta}(x)=\theta^{\top}x$. Nguyen and Ho [2022] proposal to amortize eq. 3.34 over mini-batches ${\mathcal{D}}$ of size $m$ sampled from each measure.
@@ 3.6 Policy learning for control and reinforcement learning
Many control and reinforcement learning methods amortize the solutions to a control optimization problem as illustrated in figs. 1.2 and 3.2.
Distinction. This section is on amortization for reinforcement learning and control and not the opposite direction of using reinforcement learning for amortization that section 2.2.3 discusses for parameter learning.
@@ 3.6.1 Background
A Markov decision process (MDP) can be represented with :${\mathcal{M}}\vcentcolon=({\mathcal{X}},{\mathcal{U}},p,r)$, where ${\mathcal{X}}$ are the continuous states, ${\mathcal{U}}$ are the continuous actions or controls, $p(x^{\prime}\mid x,u)$ is the transition dynamics (also referred to as the system, model, or world model), which is a Markov kernel providing the probability the next state $x^{\prime}$ given the current state-action $(x,u)$, and $r:{\mathcal{X}}\times{\mathcal{U}}$ is the reward function.
@@
which is the form that can be used to interpret many control and reinforcement learning methods as amortized optimization. Instead of amortizing the solution to eq. 3.37, methods such as Lowrey et al. [2019], Ryu et al. [2020] explicitly solve the max-Q problem.
@@
where the cost $C$ at each time is usually the negated reward $C_{t}(x_{t},u_{t})\vcentcolon=-r(x_{t},u_{t})$. The field of optimal control studies methods for solving control optimization problems of the form eq. 3.38 and standard references include Bertsekas [2000], Kirk [2004], Camacho and Alba [2013]. Eq. 3.38 induces the policy $\pi(x)\vcentcolon=u^{\star}_{1}(x)$ that solves the MDP if the horizon $H$ is long enough. Using a terminal cost at the end of the horizon can also give the controller information about how the system will perform beyond the finite-horizon rollouts being used, for example with $C_{H}(x_{H},u_{H})\vcentcolon=-Q^{\pi}(x_{H},u_{H})$.
@@
Extensions in stochastic control. The max-Q problem in eq. 3.37 can be extended to the stochastic optimization settings section 2.3.1 briefly covered when the policy $\pi$ represents a distribution over the action space ${\mathcal{U}}$. The most common objectives for stochastic policies are 1) the expected $Q$-value under the policy with
@@
where ${\mathcal{Q}}(x,\cdot)\propto\exp\left\{Q(x,\cdot)/\alpha\right\}$ is a $Q$-distribution induced by the $Q$ values that is inversely scaled by $\alpha\in\mathbb{R}_{+}$. The policy $\pi$ is usually represented as the parameters of a distribution and thus $\Pi$ becomes the space of these parameters. In most cases, $\pi$ is a Gaussian ${\mathcal{N}}(\mu,\Sigma)$ with a diagonal covariance $\Sigma$ and thus eqs. 3.39 and 3.40 can be turned into unconstrained continuous optimization problems of the form eq. 1.1 by projecting onto the Gaussian parameters. Stochastic value gradient methods such as Heess et al. [2015] often amortize eq. 3.39, while Levine and Koltun [2013], Haarnoja et al. [2018] propose methods that build on eq. 3.40, often adding additional softening and entropic terms to encourage the policy and value estimate to explore more and not converge too rapidly to a suboptimal minima. One last note is that the smoothing that the policy performs in eq. 3.39 is nearly identical to the objective smoothing considered in section 2.2.3, except in that setting the variance of the smoother remains fixed.
Connecting stochastic control back to deterministic control. This portion shows that taking stochastic policies to be Dirac delta distributions in eqs. 3.39 and 3.40 recovers the solutions to the deterministic control problem in eq. 3.37. Taking a larger classes of policy distributions, such as Gaussians, can then be interpreted as smoothing the $Q$ values to avoid 1) falling into poor local optima and 2) unstable regions where only a few actions have high value, but the rest have poor values. For the following, let $\delta_{u}(\cdot)$ be Dirac delta distribution with a parameter $u\in\mathbb{R}^{n}$ indicating the location of the mass.
@@
Let $\pi$ be a Dirac delta distribution $\delta_{u}(\cdot)$. Then the solution $\pi^{\star}(x)$ to the expected $Q$ problem in eq. 3.39 is the solution to the deterministic max-$Q$ problem in eq. 3.37.
Proof  Let $\Pi=\mathbb{R}^{n}$ be the parameter space of $\pi$ and transform eq. 3.39 to optimize over it:
@@
and thus eq. 3.41 which is the max-$Q$ operation in eq. 3.37.
Similarly for the for the KL problem in eq. 3.40:
@@
Let $\pi$ be a Dirac delta distribution $\delta_{u}(\cdot)$. Then the solution $\pi^{\star}(x)$ to the KL problem in eq. 3.40 is the solution to the deterministic max-$Q$ problem in eq. 3.37.
Proof  Let $\Pi=\mathbb{R}^{n}$ be the parameter space of $\pi$ and transform eq. 3.40 to optimize over it:
@@
${\rm D}_{\rm KL}(\delta_{u}(\cdot)\;||\;{\mathcal{Q}}(x,\cdot))$ $=\operatorname*{\mathbb{E}}_{\tilde{u}\sim\delta_{u}(\cdot)}\left[\log{\delta_{u}(\tilde{u})}-\log{\mathcal{Q}}(x,\tilde{u})\right]$ (3.44)
@@
where $C$ is a constant that does not depend on $u$, ${\mathcal{Q}}(x,u)$ is the density at $u$, and ${\mathcal{Z}}_{x}$ is the normalization constant for ${\mathcal{Q}}(x,\cdot)$ that does not depend on $u$. Putting eq. 3.44 back into eq. 3.43 and removing the constants that do not depend on $u$ gives
@@ 3.6.2 Behavioral cloning and imitation learning
This section starts in the setting where regression-based amortization is done to predict the solution of a controller that solves eq. 3.35. These settings assume access to a controller, or samples from it, that uses traditional methods, and not learning, to solve eq. 3.35 with the true or approximated dynamics. These solutions are typically available as samples from a policy $\pi^{\star}(x)$ that provides the solution to the max-Q problem in eq. 3.37 for regression-based amortization. In some settings these methods also use the optimal finite-horizon sequence $u^{\star}_{1:H}(x)$ from a solution to eq. 3.38.
@@
where, for example, the model $\pi_{\theta}$ could be a neural network. Thus BC in this setting performs regression-based amortization ${\mathcal{A}}_{\rm BC}\vcentcolon=(-Q,{\mathcal{U}},{\mathcal{X}},p(x),\pi_{\theta},{\mathcal{L}}_{\rm reg})$. Extensions from this setting could include when 1) the model $\pi_{\theta}$ is a sequence model that predicts the entire sequence $u^{\star}_{1:H}$, and 2) the MDP or policy is stochastic and eq. 3.46 turns into an amortized maximum likelihood problem rather than a regression problem.

### v5

@@ 3.4.2 Neural acceleration for matrix factorization (NeuralNMF)
where the basis matrix $W\in\mathbb{R}^{m\times r}$ and mixture matrix $H\in\mathbb{R}^{n\times r}$ are elementwise non-negative matrices of rank $r\leq\min(m,n)$. Let $Z=(W,H)$. Taking the norm of the residual of eq.˜3.22 leads to the optimization formulation
@@ 3.4.3 HyperAnderson acceleration and deep equilibrium models (HyperDEQ)
[…]
@@ 3.4.4 Comparing NeuralFP and HyperAA
$\tilde{z}^{t+1}$ $=z^{t}+{\rm diag}(\rho^{t})^{-1}(v^{t+1}-y^{t})$ (3.29)
$z^{t+1}$ $=\Pi\left(\tilde{z}^{t+1}+{\rm diag}(\rho^{t})^{-1}y^{t}\right)$
$y^{t+1}$ $=x^{t}+{\rm diag}(\rho)\left(\tilde{z}^{t+1}-z{t+1}\right),$
@@
RLQP proposes to use these OSQP iterates as a semi-amortized model with the iterates $\{s^{t},\rho^{t}\}$. They propose to only parameterize and learn to predict the step size $\rho^{t+1}\vcentcolon=\pi_{\theta}(s^{t})$, with a neural network amortization model $\pi_{\theta}$. They model the process of predicting the optimal $\rho$ as an MDP and define a reward $R_{\rm RLQP}(s,\rho)$ that is 1-1 if the QP is not solved (based on thresholds of the primal and dual residuals) and 0 otherwise, i.e. each episode rolls out the OSQP iterations with a policy predicting the optimal step size. They solve this MDP with TD3 by Fujimoto et al. [2018] to find the parameters $\theta$.
@@ 3.5 Optimal transport
where the coupling $\pi$ is joint distributions over the product of the measures, $\mathcal{U}$} is the set of admissible couplings, $c$ is a cost. The dual of eq.˜3.30, e.g. in Peyré et al. [2019, Eq. 2.31], can be represented by
@@
is the $c$-transform operation. In Euclidean spaces with the negative inner product cost [Villani, 2009, eq. 5.12], $f$ is a convex function and the $c$-transform operation $f^{c}$ in eq.˜3.33 is the standard Legendre-Fenchel transform, also known as the convex conjugate. When the measures $\alpha,\beta$ are discrete, a coupling $\pi$ in the primal (eq.˜3.30) can be represented as a matrix, and the potential $f$ in the dual (eq.˜3.31) can be represented as a finite-dimensional vector and solved with a linear programming formulation or Sinkhorn iterations [Cuturi, 2013] in the entropic setting.
The following sections overview methods that amortize these optimization problems arising for optimal transport. Section˜3.5.1 discusses methods that amortize multiple OT problems and map from the measures and cost to the optimal coupling in eq.˜3.30 or duals in eq.˜3.31. Section˜3.5.2 discusses methods that amortize the $c$-transform operation in eq.˜3.33 that arises as a repeatedly-solved subproblem when solving a single OT problem. Section˜3.5.3 overviews amortizing a slicing optimization problem that arises when projecting measures down to a single dimension for more computationally efficient solves.
@@ 3.5.1 Amortizing solutions to optimal transport problems
Many computational OT methods focus on computing the solution mapping from the measures and cost to the optimal coupling (eq.˜3.30) or duals (eq.˜3.31). When repeatedly solving optimal transport problems, this solution mapping is an optimization problem that can be amortized. These methods for predicting the optimal duals of OT problems are also related to Dinitz et al. [2021], Khodak et al. [2022], which predicts the dual solutions to matching problems. They are also related to other heuristic-based initializations for OT problems that do not use amortization such as Thornton and Cuturi [2022].
@@ 3.5.2 Amortized convex conjugates (AmorConj) and $c$-transforms
Most methods for computing the dual in eq.˜3.31 between a single pair of measures needs to repeatedly compute the $c$-transform in eq.˜3.33 to evaluate the objective value $J$ in eq.˜3.32. This transform is typically not a computational bottleneck in discrete settings when the measures have a few thousand points, e.g. in Sinkhorn solvers such as [Cuturi, 2013]. Otherwise in some continuous settings, the conjugate operation may be computationally challenging because it is an optimization problem that needs to be repeatedly solved from scratch to obtain a single Monte-Carlo estimate of $\int_{\mathcal{Y}}f^{c}(y){\rm d}\beta(y)$ to evaluate $J$ in eq.˜3.32 once.
Scoping to computing the optimal transport maps between Euclidean measures with the negative inner product cost [Villani, 2009, eq. 5.12], $f$ is a convex function and the $c$-transform operation $f^{c}$ in eq.˜3.33 is the standard Legendre-Fenchel transform. Taghvaei and Jalali [2019] discusses a lot of the theoretical foundations underpinning Wasserstein-2 optimal transport and experimentally use a numerical method that solves each conjugate operation from scratch. To alleviate the computational bottleneck of this, many methods using similar theoretical foundations amortize this conjugate operation [Dam et al., 2019, Makkuva et al., 2020, Korotin et al., 2021, Amos, 2023]. The simplest instantiation of this amortization uses a fully-amortized model $\hat{x}_{\theta}(y)$ trained with objective-based learning of the conjugate objective. Amos [2023] further discusses the modeling and loss choices in this setting and also discusses the idea of fine-tuning the amortized prediction with a numerical solver to ensure the dual objective is accurately estimated.
@@ 3.5.3 Amortized Sliced Wasserstein ($\mathcal{A}$}-SW) by Nguyen and Ho [2022]
$\textnormal{Max-SW}(\alpha,\beta)\vcentcolon=\max_{\theta\in{\mathcal{S}}^{d-1}}W(\theta;\alpha,\beta)\qquad W(\theta;\alpha,\beta)\vcentcolon=W(\theta_{\#}\alpha,\theta_{\#}\beta),$ (3.34)
where $\theta_{\#}\alpha$ is the push-forward measure of $\mu$ through $T_{\theta}(x)=\theta^{\top}x$. Nguyen and Ho [2022] propose to amortize eq.˜3.34 over mini-batches $\mathcal{D}$} of size $m$ sampled from each measure.
@@ 3.6 Policy learning for control and reinforcement learning
Many control and reinforcement learning methods amortize the solutions to a control optimization problem as illustrated in figs.˜1.2 and 3.2.
Distinction. This section is on amortization for reinforcement learning and control and not the opposite direction of using reinforcement learning for amortization that section˜2.2.3 discusses for parameter learning.
@@ 3.6.1 Background
A Markov decision process (MDP) can be represented with ${\mathcal{M}}\vcentcolon=({\mathcal{X}},{\mathcal{U}},p,r)$, where $\mathcal{X}$} are the continuous states, $\mathcal{U}$} are the continuous actions or controls, $p(x^{\prime}\mid x,u)$ is the transition dynamics (also referred to as the system, model, or world model), which is a Markov kernel providing the probability the next state $x^{\prime}$ given the current state-action $(x,u)$, and $r:{\mathcal{X}}\times{\mathcal{U}}$ is the reward function.
@@
which is the form that can be used to interpret many control and reinforcement learning methods as amortized optimization. Instead of amortizing the solution to eq.˜3.37, methods such as Lowrey et al. [2019], Ryu et al. [2020] explicitly solve the max-Q problem.
@@
where the cost $C$ at each time is usually the negated reward $C_{t}(x_{t},u_{t})\vcentcolon=-r(x_{t},u_{t})$. The field of optimal control studies methods for solving control optimization problems of the form eq.˜3.38 and standard references include Bertsekas [2000], Kirk [2004], Camacho and Alba [2013]. Eq.˜3.38 induces the policy $\pi(x)\vcentcolon=u^{\star}_{1}(x)$ that solves the MDP if the horizon $H$ is long enough. Using a terminal cost at the end of the horizon can also give the controller information about how the system will perform beyond the finite-horizon rollouts being used, for example with $C_{H}(x_{H},u_{H})\vcentcolon=-Q^{\pi}(x_{H},u_{H})$.
@@
Extensions in stochastic control. The max-Q problem in eq.˜3.37 can be extended to the stochastic optimization settings section˜2.3.1 briefly covered when the policy $\pi$ represents a distribution over the action space $\mathcal{U}$}. The most common objectives for stochastic policies are 1) the expected $Q$-value under the policy with
@@
where ${\mathcal{Q}}(x,\cdot)\propto\exp\left\{Q(x,\cdot)/\alpha\right\}$ is a $Q$-distribution induced by the $Q$ values that is inversely scaled by $\alpha\in\mathbb{R}_{+}$. The policy $\pi$ is usually represented as the parameters of a distribution and thus $\Pi$ becomes the space of these parameters. In most cases, $\pi$ is a Gaussian ${\mathcal{N}}(\mu,\Sigma)$ with a diagonal covariance $\Sigma$ and thus eqs.˜3.39 and 3.40 can be turned into unconstrained continuous optimization problems of the form eq.˜1.1 by projecting onto the Gaussian parameters. Stochastic value gradient methods such as Heess et al. [2015] often amortize eq.˜3.39, while Levine and Koltun [2013], Haarnoja et al. [2018] propose methods that build on eq.˜3.40, often adding additional softening and entropic terms to encourage the policy and value estimate to explore more and not converge too rapidly to a suboptimal minima. One last note is that the smoothing that the policy performs in eq.˜3.39 is nearly identical to the objective smoothing considered in section˜2.2.3, except in that setting the variance of the smoother remains fixed.
Connecting stochastic control back to deterministic control. This portion shows that taking stochastic policies to be Dirac delta distributions in eqs.˜3.39 and 3.40 recovers the solutions to the deterministic control problem in eq.˜3.37. Taking a larger classes of policy distributions, such as Gaussians, can then be interpreted as smoothing the $Q$ values to avoid 1) falling into poor local optima and 2) unstable regions where only a few actions have high value, but the rest have poor values. For the following, let $\delta_{u}(\cdot)$ be Dirac delta distribution with a parameter $u\in\mathbb{R}^{n}$ indicating the location of the mass.
@@
Let $\pi$ be a Dirac delta distribution $\delta_{u}(\cdot)$. Then the solution $\pi^{\star}(x)$ to the expected $Q$ problem in eq.˜3.39 is the solution to the deterministic max-$Q$ problem in eq.˜3.37.
Proof Let $\Pi=\mathbb{R}^{n}$ be the parameter space of $\pi$ and transform eq.˜3.39 to optimize over it:
@@
and thus eq.˜3.41 which is the max-$Q$ operation in eq.˜3.37.
Similarly for the for the KL problem in eq.˜3.40:
@@
Let $\pi$ be a Dirac delta distribution $\delta_{u}(\cdot)$. Then the solution $\pi^{\star}(x)$ to the KL problem in eq.˜3.40 is the solution to the deterministic max-$Q$ problem in eq.˜3.37.
Proof Let $\Pi=\mathbb{R}^{n}$ be the parameter space of $\pi$ and transform eq.˜3.40 to optimize over it:
@@
[…]
@@
where $C$ is a constant that does not depend on $u$, ${\mathcal{Q}}(x,u)$ is the density at $u$, and ${\mathcal{Z}}_{x}$ is the normalization constant for ${\mathcal{Q}}(x,\cdot)$ that does not depend on $u$. Putting eq.˜3.44 back into eq.˜3.43 and removing the constants that do not depend on $u$ gives
@@ 3.6.2 Behavioral cloning and imitation learning
This section starts in the setting where regression-based amortization is done to predict the solution of a controller that solves eq.˜3.35. These settings assume access to a controller, or samples from it, that uses traditional methods, and not learning, to solve eq.˜3.35 with the true or approximated dynamics. These solutions are typically available as samples from a policy $\pi^{\star}(x)$ that provides the solution to the max-Q problem in eq.˜3.37 for regression-based amortization. In some settings these methods also use the optimal finite-horizon sequence $u^{\star}_{1:H}(x)$ from a solution to eq.˜3.38.
@@
where, for example, the model $\pi_{\theta}$ could be a neural network. Thus BC in this setting performs regression-based amortization ${\mathcal{A}}_{\rm BC}\vcentcolon=(-Q,{\mathcal{U}},{\mathcal{X}},p(x),\pi_{\theta},{\mathcal{L}}_{\rm reg})$. Extensions from this setting could include when 1) the model $\pi_{\theta}$ is a sequence model that predicts the entire sequence $u^{\star}_{1:H}$, and 2) the MDP or policy is stochastic and eq.˜3.46 turns into an amortized maximum likelihood problem rather than a regression problem.